How did you go from .055(m(solute)+600))=m(solute) to -0.945•m(solute)=-33
1
Expert's answer
2019-05-01T08:01:06-0400
"0.055(m(solute)+600)= m(solute)"

"0.055\times m(solute)+ 0.055\times 600= m(solute)"

"0.055\times m(solute)+33= m(solute)"

"0.055\times m(solute)-m(solute)= -33"

"m(solute)\times(0.055-1)=-33"

"-0.945\times m(solute)=-33"

"m(solute)=\frac{-33}{-0.945}"

"m(solute)= 34.92"